# Flow-Inspired Lightweight Multi-Robot Real-Time Scheduling Planner

### v1

@@ Abstract
Collision avoidance and trajectory planning are crucial in multi-robot systems, particularly in environments with numerous obstacles. Although extensive research has been conducted in this field, the challenge of rapid traversal through such environments has not been fully addressed. This paper addresses this problem by proposing a novel real-time scheduling scheme designed to optimize the passage of multi-robot systems through complex, obstacle-rich maps. Inspired from network flow optimization, our scheme decomposes the environment into a network structure, enabling the efficient allocation of robots to paths based on real-time congestion data. The proposed scheduling planner operates on top of existing collision avoidance algorithms, focusing on minimizing traversal time by balancing robot detours and waiting times. Our simulation results demonstrate the efficiency of the proposed scheme. Additionally, we validated its effectiveness through real world flight tests using ten quadrotors. This work contributes a lightweight, effective scheduling planner capable of meeting the real-time demands of multi-robot systems in obstacle-rich environments.
@@ I INTRODUCTION
In obstacle-rich environments, the primary challenge for multi-robot navigation is effective trajectory planning. Methods such as the artificial potential field method [1] [2], velocity obstacle (VO) and its variants [3] [4] [5], on-demand collision avoidance [6], gradient-based local planning methods [7], and distributed model predictive control (DMPC) [8] have been proposed to address it. These approaches mainly focus on collision avoidance, reduce computational load, minimize deadlocks, and handle complex environments. However, few methods focus on enabling rapid traversal through these areas.
@@
However, real-time and efficient scheduling of multi-robot systems in obstacle-rich environments poses significant challenges. Flow-based methods cannot be directly applied to multi-robot systems for several reasons. First, flow-based methods require a predefined network structure, but real-world map environments often lack such readily usable network structures. Second, these methods typically have long computation times and are often designed for offline use, lacking real-time applicability. Additionally, flow-based methods do not account for robot collision avoidance and motion models. In scenarios with a large number of robots, collision avoidance becomes a primary cause of congestion. At this point, robots face a decision: either wait in congested areas or choose a longer but uncongested path. Therefore, a new scheduling scheme is required, one that can determine in real time which robots should wait and which should reroute, while also specifying alternative routes. Moreover, the scheduling process needs to allocate the number of robots based on path capacity to reduce congestion.
To address these challenges, we propose a real-time scheduling scheme for multi-robot systems. Our planner builds on existing collision avoidance algorithms. It does not directly solve robot collision avoidance but performs real-time scheduling based on the map environment and robot status. The scheduling scheme first decomposes the initial map and extracts information to form a network containing nodes, edges, and capacity information. Subsequently, the planner calculates the path for each robot in real-time based on the current position of the robots and the congestion level of the map channels. The planner optimizes path selection from the perspective of the entire swarm, balancing detouring and waiting to minimize the time required for robots to traverse obstacle-rich environments. Experimental results show that when the number of robots reaches 500, the calculation time per instance of our planner is only 0.9 seconds, demonstrating high computational efficiency. And we conducted flight tests to demonstrate that the proposed scheme can be applied to real-world settings.
@@
  • We propose a novel scheduling planner that utilizes the traversable areas of the map, thereby reducing the traversal time for a swarm of robots in obstacle-rich environments.
  • Our scheduling planner is lightweight and exhibits high computational efficiency, meeting the real-time requirements of the robots.
  • We have validated the effectiveness of our algorithm through extensive experiments.
@@
Many methods have been developed to address the navigation challenges of multiple robots in obstacle-rich environments, each with a distinct focus. The artificial potential field method is often favored for its low computational load and high scalability [1], [2]. Velocity obstacle (VO) based methods and their variants have shown good performance in multi-robot navigation scenarios [3], [4], [5]. Optimization-based approaches, including on-demand collision avoidance [6], gradient-based local planning [7], and distributed model predictive control (DMPC) [8] aim to reduce computational time. Alonso-Mora et al. [10] introduced a planning method for multi-robot task execution in dynamic environments, ensuring task specification compliance and collision avoidance. Tordesillas et al. [11] proposed a decentralized and asynchronous trajectory planner that offers faster, safer, and more efficient navigation. Park et al. [12] presented a trajectory planning algorithm that uses dynamic linear safety corridors to construct feasible collision constraints, enabling deadlock-free planning in extreme environments. While these studies primarily focus on collision avoidance, reducing computational load, and managing complex scenarios, there is limited research on minimizing traversal time for multi-robot systems in obstacle-rich environments.
Since the pioneering research of Ford and Fulkerson on the maximum flow problem in networks [9], this field has attracted considerable attention. Variants such as the multi-commodity flow problem share similarities with multi-robot navigation challenges. The multi-commodity flow problem aims to transport a set of commodities from sources to destinations while respecting network capacity constraints and minimizing overall transportation cost or time. Trivella et al. [13] proposed an algorithm that balances delays and costs using penalty functions, enhancing the flexibility and cost efficiency of liner shipping networks. Bevrani et al. [14] introduced a multi-criteria multi-commodity network flow model to evaluate the resilience and vulnerability of transportation systems. Huang et al. [15] proposed a game-theoretic model and algorithm to optimize the decision-making process of multiple vehicles in a network. Khanal et al. [16] addressed the maximum static and dynamic multicommodity flow problem with intermediate storage. These methods generally require prior knowledge of network information and focus on adhering to capacity limits or reducing transportation costs. Their computational time is often lengthy, making them unsuitable for direct application in multi-robot navigation.
@@ III Problem Statement
To fulfill the mission requirements, robots must navigate from starting points, traverse an obstacle-rich environment, and ultimately reach their destinations. During this process, robots must account for collision avoidance both among themselves and with the obstacles. The efficiency of the overall traversal of multi-robot system is measured by the time taken for the last robot to reach the destination. In scenarios where path capacity is limited and the number of robots is large, collision avoidance can lead to congestion, thereby decreasing the overall traversal efficiency.
To address this issue, we need to consider: (1) How to allocate the number of robots based on path capacity to reduce congestion. (2) When facing congestion ahead, whether a robot should wait or reroute, and if rerouting, which path to select. (3) How to make decisions between a shorter but congested path and a longer but clear path. As our scheme focuses on scheduling, we will adopt existing solutions for collision avoidance algorithms. In the experimental section, we will employ the Reciprocal Velocity Obstacles (RVO) algorithm [3] for collision avoidance.
@@ IV Network construction
In this section, we delve into the details of network construction. The initial map is typically a grid map containing information on obstacles and traversable areas. To construct the subsequent network, we first decompose the map using the Boustrophedon Cellular Decomposition method [17]. This method scans the map by moving back and forth, similar to the way oxen plow fields, dividing the map into multiple cells. Denote the cell set containing all cells as $\mathcal{C}$. Next, we will introduce how to construct “$\operatorname{PathNode}$” and “$\operatorname{PathPos}$” at the boundary position of the cells.
Figure 1: The network construction process. The blue circles represent the “$\operatorname{PathNode}$” of the network, and the red circles represent the “$\operatorname{PathPos}$” of each node.
IV-A $\operatorname{PathNode}$” and “$\operatorname{PathPos}$
We extract the length and positional information of each boundary of cells and establish passage positions at the shared boundaries of adjacent cells. Denote the shared boundary length as $L_{B}$. Then, the number of passage positions that can be established on the boundary is $\frac{L_{B}}{\alpha\;r_{\min}}$, where $\alpha$ is a constant. Typically $\alpha$ is set to $1\sim 2$, to ensure sufficient maneuvering space for the robots. The number of passage positions represents the capacity of boundary, which is the maximum number of robots that can pass through simultaneously. These passage positions are defined as “$\operatorname{PathPos}$” (red circles in Fig.(1)). Next, we group every $N_{B}$ passage positions together, partitioning the red positions into multiple groups. Then, we place a new node at the mean position of each group. The new node is defined as “$\operatorname{PathNode}$” ( blue circles in Fig.(1)). If the remaining number of red positions is less than $N_{B}$, they form a separate group. “$\operatorname{PathNode}$” and “$\operatorname{PathPos}$” serve as bridges between the actual map and the network graph for the robots.
@@ V Scheduling Planner
In this section, we discuss the details of our scheduling planner. As illustrated in Fig.(4), our scheduling planner consists of three main components: path set search, path node selection and local position allocation. The path node selection component calculates $\operatorname{PathNode}^{(n)}(t)$ for each robot, while the local position allocation component determines the corresponding $\operatorname{PathPos}^{(n)}(t)$. After these computations, the control variables are transmitted to the swarm robots, guiding them along their designated paths.
@@ V-A Path set search
Define the next-go-to path set for the $n$th robot as $\mathcal{P}^{(n)}_{\text{next}}(t)$, which represents the set of paths that the robot can potentially take from its current state in the future. During each decision-making instance, the robot considers only the paths within this set. The path search process is as follows: (1) Identify all $\operatorname{PathNode}$s on the upper boundary of the cell where the robot is currently located $\operatorname{UP}(c^{(n)}_{now}(t))$; (2) Identify all $\operatorname{PathNode}$s on the lower boundary of the goal cell $\operatorname{DN}(c_{goal}^{(n)})$; (3) Compute the paths from all upper boundary $\operatorname{PathNode}$s to the lower boundary $\operatorname{PathNode}$s using Dijkstra’s algorithm. (4) Add the resulting paths to path set. The next-go-to path set $\mathcal{P}^{(n)}_{\text{next}}(t)$ can be written as:
@@
Thus, we find the next-go-to path set for the $n$th robot. The rationale behind this method of computing the path set is that our algorithm is designed for real-time operation, rather than offline processing. Therefore, it is unnecessary to spend excessive time finding all possible intermediate paths. Instead, we only need to consider the viable paths from the robot’s current state to the destination. Our experiments have demonstrated that this search method is highly efficient, with the time for a single search involving 500 robots being approximately 0.4 seconds, and the final decision-making performance is satisfactory.
@@ V-B1 Link queuing cost
$\sum\limits_{l\in\operatorname{Fir}(\mathcal{P}_{\text{next}}(t))}\!\!\!\!\!\frac{\left(\sum\limits_{n\in\mathcal{N}}\sum\limits_{p\in\mathcal{P}^{l}_{\text{fir}}(t)}z_{p}^{(n)}(t)+\operatorname{Num}(l)-\operatorname{Cap}(l)\right)^{2}}{\operatorname{Cap}(l)^{2}}$ (14)
The term $\sum_{n\in\mathcal{N}}\sum_{p\in\mathcal{P}^{l}_{\text{fir}}(t)}z_{p}^{(n)}(t)$ represents the number of times link $l$ is selected across all robots path results. When added to the current number of robots on link $l$, this gives the total number of robots on link $l$ after decision-making. We then subtract the capacity of link $l$, and compute the square of the resulting value. This represents the queuing cost for the first link. Summing the queuing cost for all first links yields $f_{\operatorname{Fir}}(t)$. This cost term ensures that the path node selection does not exceed the capacity of the links. When there are already many robots on an link, the remaining capacity is limited, and assigning more robots to that link will increase the loss. Similarly, if the sum of the assigned and existing robots is less than the capacity, the loss will also increase, encouraging the decision-making process to fully utilize the link capacities.
Define the queuing cost of the second link as $f_{\operatorname{Sec}}(t)$,
$\sum\limits_{l\in\operatorname{Sec}(\mathcal{P}_{\text{next}}(t))}\!\!\!\!\!\!\frac{\left(\sum\limits_{n\in\mathcal{N}}\sum\limits_{p\in\mathcal{P}^{l}_{\text{sec}}(t)}z_{p}^{(n)}(t)+\operatorname{Num}(l)-\operatorname{Cap}(l)\right)^{2}}{\operatorname{Cap}(l)^{2}}$ (15)
@@ V-C Local position allocation
Upon obtaining the path node selection results $\operatorname{PathNode}^{(n)}(t)$, the next step is to allocate specific node positions $\operatorname{PathPos}^{(n)}(t)$ for the robots. Each node has multiple positions, and we utilize a greedy algorithm for this allocation. Initially, we select the position closest to the current position of robot. For each subsequent node, we select the position that is nearest to the previously allocated position. This process is repeated iteratively, ensuring each chosen position is the closest to the last assigned position. We employ a greedy algorithm instead of more complex alternatives. Our experiments show that the greedy algorithm is both simple and efficient, with the allocation time for 500 robots being approximately 0.02 seconds.

### v2

# Flow-Inspired Multi-Robot Real-Time Scheduling Planner

@@ Abstract
Collision avoidance and trajectory planning are crucial in multi-robot systems, particularly in environments with numerous obstacles. Although extensive research has been conducted in this field, the challenge of rapid traversal through such environments has not been fully addressed. This paper addresses this problem by proposing a novel real-time scheduling scheme designed to optimize the passage of multi-robot systems through complex, obstacle-rich maps. Inspired from network flow optimization, our scheme decomposes the environment into a network structure, enabling the efficient allocation of robots to paths based on real-time congestion data. The proposed scheduling planner operates on top of existing collision avoidance algorithms, focusing on minimizing traversal time by balancing robot detours and waiting times. Our simulation results demonstrate the efficiency of the proposed scheme. Additionally, we validated its effectiveness through real world flight tests using ten drones. This work contributes an effective scheduling planner capable of meeting the real-time demands of multi-robot systems in obstacle-rich environments.
@@ I INTRODUCTION
In obstacle-rich environments, the primary challenge for multi-robot navigation is effective trajectory planning. Methods such as the velocity obstacle [1] [2], distributed model predictive control [3] [4], gradient-based local planning methods [5] and optimization-based methods [6] [7] have been proposed to address it. These approaches typically employ a two-stage process: front-end path planning for discrete path search and back-end collision avoidance optimization to refine trajectories. Common front-end path planning methods include point-to-point [1, 2, 3, 4], Astar [5], and methods like ECBS (Enhanced Conflict Based Search) [6] [7]. These studies mainly focus on the back-end, dealing with collision avoidance, reducing computational load, and minimizing deadlock. However, we find that there is still much room for improvement in the front-end path planning, especially in the case where multiple robots need to rapidly traverse an environment with obstacles.
@@
However, real-time and efficient scheduling of multi-robot systems in obstacle-rich environments poses significant challenges. Flow-based methods cannot be directly applied to multi-robot systems for several reasons. First, flow-based methods require a predefined network structure, but real-world map environments often lack such readily usable network structures. Second, these methods typically have long computation times and are often designed for offline use, lacking real-time applicability. Additionally, flow-based methods do not account for robot collision avoidance and motion models. In scenarios with a large number of robots, collision avoidance becomes a primary cause of congestion. At this point, robots face a decision: either wait in congested areas or choose a longer but uncongested path. Therefore, a new scheduling scheme is required, one that can determine in real time which robots should wait and which should reroute, while also specifying alternative routes.
To address these challenges, we propose a real-time scheduling scheme for multi-robot systems. Our planner builds on existing collision avoidance algorithms. It does not directly solve robot collision avoidance but performs real-time scheduling based on the map environment and robot status. The scheduling scheme first decomposes the initial map and extracts information to form a network containing nodes, edges, and capacity information. Subsequently, the planner calculates the path for each robot in real-time based on the current position of the robots and the congestion level of the map channels. The planner optimizes path selection from the perspective of the entire swarm, balancing detouring and waiting to minimize the time required for robots to traverse obstacle-rich environments. Experimental results show that our planner has high computational efficiency. And we conducted flight tests to demonstrate that the proposed scheme can be applied to real-world settings.
@@
  • We propose a scheduling planner that utilizes the traversable areas of the map, thereby reducing the traversal time for multi-robots in obstacle-rich environments.
  • Our scheduling planner exhibits high computational efficiency, meeting the real-time requirements of the robots. And we have validated the effectiveness of our algorithm through extensive experiments.
  • To the best of our knowledge, we are the first to apply the flow-based idea to multi-robot real-time scheduling, in order to improve the efficiency of traversing environments with multiple obstacles.
@@
Numerous methods have been devised to tackle the navigation challenges of multiple robots in environments filled with obstacles, each with its own emphasis. Methods based on Velocity Obstacle (VO) and their derivatives have exhibited good performance in multi-robot navigation scenarios [1][2]. Luis et al.[3] introduced a distributed model predictive control algorithm for real-time trajectory generation of multiple robots. Soria et al.[4] presented a predictive model that optimizes potential field principles to improve the speed, order, and safety of aerial swarms in complex environments. The front-end paths in these studies [1, 2, 3, 4] are rather straightforward, typically being point-to-point paths. Zhou et al.[5] proposed a gradient-based local planning method for multi-robot systems, where the front-end path search uses Astar. Park et al.[6] proposed a method that combines ECBS and bernstein polynomial-based optimization for planning. Li et al.[7] proposed a planning method that merges ECBS with a priority-based nonlinear optimization approach. Both [6] and [7] use ECBS as a multi-agent path finding method for front-end search. Although these studies mainly concentrate on collision avoidance, reducing computational load, and handling complex scenarios, they overlook the time delays caused by congestion as the number of robots increases. As a result, when the map becomes more intricate and the number of robots grows, the efficiency of these front-end search methods drops significantly.
Since Ford and Fulkerson’s pioneering work on the maximum flow problem in networks [8], this area has received substantial attention. Variations like the multi-commodity flow problem share similarities with multi-robot navigation challenges. The multi-commodity flow problem aims to transport a set of commodities from sources to destinations while adhering to network capacity constraints and minimizing the overall transportation cost or time [9, 10, 11, 12]. However, these methods generally require prior knowledge of network information and focus on strictly meeting network capacity and boundary constraints. Their computational time is often long, rendering them unsuitable for direct application in multi-robot navigation. There are also some studies attempting to apply flow-based concepts to robot path planning. Yu et al.[13] showed that multi-agent path planning can be reduced to a network flow problem and proposed an algorithm with a complexity of $O(nVE)$. Dugas et al.[14] modeled crowds as a flow model and proposed a method for robot navigation within crowds. Janchiv et al.[15] applied flow network techniques to the multi-robot complete coverage path planning problem. Nevertheless, no research has yet applied flow-based ideas to multi-robot real-time scheduling and planning.
@@ III Problem Statement
To fulfill the mission requirements, robots must navigate from starting points, traverse an obstacle-rich environment, and ultimately reach their destinations. During this process, robots must account for collision avoidance both among themselves and with the obstacles. The efficiency of the overall traversal of multi-robot system is measured by the time taken for the last robot to reach the destination.
Figure 1: A motivation example.
Fig.1 shows a motivation example. Ten robots need to move from the starting position to the target position. There are two selectable paths in the figure, namely Path1 and Path2. Path1 can only allow one robot to pass through each time, and it takes one second to pass. Path2 can allow two robots to pass through simultaneously each time, and it takes 3 seconds to pass. If all ten robots choose the shorter path Path1, then it will take ten seconds for all robots to reach the target position. However, if six robots are assigned to Path1 and four robots to Path2, then the six robots choosing Path1 will take six seconds to pass, and the four robots passing through Path2 simultaneously will also take six seconds. Thus, under this strategy, it only takes six seconds for all ten robots to pass through this area.
This example shows that by allocating some robots to paths with longer lengths but higher capacities, the overall efficiency can be improved. For the more complex situation (a more complex map, a larger number of robots, and the need to consider the motion model and collision avoidance). How to achieve real-time scheduling and planning, and reasonably allocate the number of robots according to the capacity of the paths to improve efficiency is not obvious. Therefore, a sophisticated scheduling scheme is needed.
@@ IV Network construction
In this section, we will thoroughly explore the details of network construction. The initial map is typically a grid map that includes information about obstacles and passable areas. The network construction process involves extracting the passable area information from the grid map and then constructing a network with path and capacity information, which will serve as the foundation for subsequent scheduling. Note that unlike the planner that runs in real-time, the process of network construction only needs to be initialized once.
To build the subsequent network, we first employ the Boustrophedon Cellular Decomposition method [16] for map decomposition. This method scans the map in a back and forth motion, much like how oxen plow fields, dividing the map into multiple cells. Let the set of all cells be denoted as $\mathcal{C}$. Through map decomposition, the entire map is partitioned into multiple distinct regions, which we term cells. Each cell is connected to others and has adjacent boundary information. After processing this boundary information, we can introduce the concepts of $\operatorname{PathNode}$ and $\operatorname{PathPos}$.
Figure 2: The network construction process. The blue circles represent the $\operatorname{PathNode}$ of the network, and the red circles represent the $\operatorname{PathPos}$ of each node.
IV-A $\operatorname{PathNode}$ and $\operatorname{PathPos}$
$\operatorname{PathNode}$ and $\operatorname{PathPos}$ refer to some key positions on the adjacent boundaries of cells, which characterize the traffic capacity of the boundaries. First, we extract the length and position information of the boundaries of each cell, and determine the traffic positions on the common boundaries of adjacent cells. Denote the length of the common boundary as $L_{B}$. Then, the number of traffic positions that can be determined on this boundary is $\frac{L_{B}}{\alpha\;r_{\min}}$, where $\alpha$ is a constant, usually taking values between 1 and 2 to ensure that the robot has sufficient maneuvering space. The number of passage positions reflects the capacity of the boundary, that is, the maximum number of robots that can pass through simultaneously. These passage positions are defined as $\operatorname{PathPos}$ (shown as red circles in Figure (2)). Next, we group the passage positions on the same boundary. Every $N_{B}$ passage positions are grouped into one group, where $N_{B}$ represents the number of positions in each group. Subsequently, a new node is placed at the average position of each group, and this new node is defined as $\operatorname{PathNode}$ (shown as blue circles in Figure (2)). If the number of remaining red positions is less than $N_{B}$, these positions form a separate group. $\operatorname{PathNode}$ and $\operatorname{PathPos}$ play an important role as the bridge between the map and the network.
@@ V Scheduling Planner
In this section, we discuss the details of our scheduling planner. As illustrated in Fig.(5), our scheduling planner consists of three main components: path set search, path node selection and local position allocation. First, path set search will search for potential paths based on the current state of the robots in the network, and place these screened out paths in an alternative set. Subsequently, path node selection will comprehensively consider the path alternative sets of all robots, taking into account factors such as path length, path capacity, and the degree of robot congestion, and select appropriate path nodes for each robot. Finally, local position allocation will select a precise path position for each robot, that is, the passage position on the boundary. Through the collaborative operation of these parts, the control variables $\operatorname{PathNode}^{(n)}(t)$ and $\operatorname{PathPos}^{(n)}(t)$ are transmitted to the swarm robots, guiding them along their designated paths.
@@ V-A Path set search
Define the next-go-to path set for the $n$th robot as $\mathcal{P}^{(n)}_{\text{next}}(t)$, which represents the set of paths that the robot can potentially take from its current state in the future. During each decision-making instance, the robot considers only the paths within this set. The path search process is as follows: (1) Identify all $\operatorname{PathNode}$s on the upper boundary of the cell where the robot is currently located $\operatorname{UP}(c^{(n)}_{now}(t))$; (2) Identify all $\operatorname{PathNode}$s on the lower boundary of the goal cell $\operatorname{DN}(c_{goal}^{(n)})$; (3) Compute the paths from all upper boundary $\operatorname{PathNode}$s to the lower boundary $\operatorname{PathNode}$s using Dijkstra algorithm; (4) Add the resulting paths to path set. The next-go-to path set $\mathcal{P}^{(n)}_{\text{next}}(t)$ can be written as:
@@
Thus, we find the next-go-to path set for the $n$th robot.
Intuitively, we can understand this search process in the following way. First, find out which exit nodes are available for selection in the cell where the current robot is located. Afterwards, determine which entrance nodes can be chosen to reach the target cell. Then, use the Dijkstra algorithm directly to find the shortest paths connecting the nodes, and add these paths to the set of selectable paths. We directly use the Dijkstra algorithm to find the shortest paths between nodes without considering the intermediate branching paths. The reason for this is that our algorithm is designed for real-time operation, rather than offline processing. Therefore, it is unnecessary to spend excessive time finding all possible intermediate paths. Instead, we only need to consider the viable paths from the robot’s current state to the destination. Our experiments have demonstrated that this search method is highly efficient, with the time for a single search involving 500 robots being approximately 0.4 seconds, and the final decision-making performance is satisfactory.
@@ V-B1 Link queuing cost
$\sum\limits_{l\in\operatorname{Fir}(\mathcal{P}_{\text{next}}(t))}\!\!\!\!\!\frac{\left(\sum\limits_{n\in\mathcal{N}}\sum\limits_{p\in\mathcal{P}^{l}_{\text{fir}}(t)}z_{p}^{(n)}(t)+\operatorname{Num}(l)-\operatorname{Cap}(l)\right)^{2}}{\operatorname{Cap}(l)^{2}}.$ (14)
The term $\sum_{n\in\mathcal{N}}\sum_{p\in\mathcal{P}^{l}_{\text{fir}}(t)}z_{p}^{(n)}(t)$ represents the number of times link $l$ is selected across all robots path results. When added to the current number of robots on link $l$, this gives the total number of robots on link $l$ after decision-making. We then subtract the capacity of link $l$, and compute the square of the resulting value. This represents the queuing cost for the first link. Summing the queuing cost for all first links yields $f_{\operatorname{Fir}}(t)$. This cost term ensures that the path node selection does not exceed the capacity of the links. When there are already many robots on an link, the remaining capacity is limited, and assigning more robots to that link will increase the loss. Similarly, if the sum of the assigned and existing robots is less than the capacity, the loss will also increase, encouraging the decision-making process to fully utilize the link capacities. Define the queuing cost of the second link as $f_{\operatorname{Sec}}(t)$:
$\sum\limits_{l\in\operatorname{Sec}(\mathcal{P}_{\text{next}}(t))}\!\!\!\!\!\!\frac{\left(\sum\limits_{n\in\mathcal{N}}\sum\limits_{p\in\mathcal{P}^{l}_{\text{sec}}(t)}z_{p}^{(n)}(t)+\operatorname{Num}(l)-\operatorname{Cap}(l)\right)^{2}}{\operatorname{Cap}(l)^{2}}.$ (15)
@@ V-C Local position allocation
Upon obtaining the path node selection results $\operatorname{PathNode}^{(n)}(t)$, the next step is to allocate specific node positions $\operatorname{PathPos}^{(n)}(t)$ for the robots. Each node has multiple positions, and we adopt a method of selecting the positions with the closest distance one by one. Initially, we select the position closest to the current position of robot. For each subsequent node, we select the position that is nearest to the previously allocated position. This process is repeated iteratively, ensuring each chosen position is the closest to the last assigned position. In this section, we directly choose the position based on the closest distance, rather than more complex methods. We found in the experiments that this assignment method works very well. When robots cannot reach the ideal positions due to congestion and choose the positions closest to themselves, the previous section “Path node selection” will replan new nodes for them, thus improving the overall efficiency. Therefore, this algorithm is simple and efficient.